40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 9919     Entry time: Tue May 6 19:38:13 2014
Author: Jenne 
Type: Update 
Category: LSC 
Subject: Set up for PRFPMI CM locking 

To get ready for the PRFPMI CM trials tonight, I put AS55's cables back to their nominal state, and now have REFL11 I going to IN1 of the CM board.  The OUT1 of the CM board goes to the REFL11I whitening channel.

REFLDC was not disconnected in the last few days, so it is still set up for IN2 of the CM board, with an external offset adjust.

ELOG V3.1.3-